OpenStudy (anonymous):

solve : ln x = 1 + ln (x+1)

OpenStudy (aravindg):

@camilasanchez where exactly are you stuck in this question?

OpenStudy (blurbendy):

ln x = 1 + ln (x+1) ln|x| -1 - ln|x + 1| = 0 combine the logs ln | (x) / (x +1) | -1 = 0 ln | (x) / (x +1) | = 1 (x) / (x + 1) = e x = e(x + 1) x = ex + e (1 - e)x = e x = (e) / (1 -e)

OpenStudy (anonymous):

ln(x)-ln(x+1)=1 ln(x/x+1)=ln(e) (x)/(x+1)=e x=(x+1)e x=xe+e x(1-e)=e x=(e)/(e-1)

Directrix (directrix):

@nitz Check your penultimate step and the last step, please. I didn't follow how the (1-e) became (e-1). Thanks.

OpenStudy (anonymous):

sorry.....its x=xe+e x-xe=e x(1-e)=e x=e/(1-e)

hartnn (hartnn):

just a reminder... don't post direct entire solutions, its against CoC.

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

sorry @hartnn